# VitaGlyph: Vitalizing Artistic Typography with Flexible Dual-branch Diffusion Models

### v2

@@ Abstract
Artistic typography is a technique to visualize the meaning of input character in an imaginable and readable manner. With powerful text-to-image diffusion models, existing methods directly design the overall geometry and texture of input character, making it challenging to ensure both creativity and legibility. In this paper, we introduce a dual-branch, training-free method called VitaGlyph, enabling flexible artistic typography with controllable geometry changes while maintaining the readability. The key insight of VitaGlyph is to treat input character as a scene composed of a Subject and its Surrounding, which are rendered with varying degrees of geometric transformation. To enhance the visual appeal and creativity of the generated artistic typography, the subject flexibly expresses the essential concept of the input character, while the surrounding enriches relevant background without altering the shape, thus maintaining overall readability. Specifically, we implement VitaGlyph through a three-phase framework: (i) Knowledge Acquisition leverages large language models to design text descriptions for the subject and surrounding. (ii) Regional Decomposition detects the part that most closely matches the subject description and divides the input glyph image into subject and surrounding regions. (iii) Typography Stylization first refines the structure of the subject region via Semantic Typography, and then separately renders the textures of the Subject and Surrounding regions through Controllable Compositional Generation. Experimental results demonstrate that VitaGlyph not only achieves better artistry and readability but also manages to depict multiple customized concepts, facilitating more creative and pleasing artistic typography generation. Our code will be made publicly available.
Figure 1: Examples of our VitaGlyph. The rendered images not only convey the intrinsic semantics of input characters (i.e., first row) but also visualize the user-specific concepts (i.e., second row).
@@ 1 Introduction
Figure 2: Overview of VitaGlyph. VitaGlyph divides the input glyph image into Subject and Surrounding and follows a three-phase pipeline: (i) Given an input character, Knowledge Acquisition leverages large language models to design prompts for both subject and surrounding. (ii) With the subject prompt and glyph image, Regional decomposition detects the part that best matches the subject description and splits the glyph into subject and surrounding regions. (iii) Typography Stylization first transforms the Subject structure via Semantic Typography and then separately renders textures for both regions through Controllable Compositional Generation.
Artistic typography focuses on emphasizing the visual meaning of words while maintaining their legibility. This technique has widespread applications across various fields such as commercial advertising [10, 22], social education [43, 9], and design promotion [52, 45]. Traditional artistic typography involves manually applying design principles such as color, spacing, and font choice to convey meaning [11, 17]. However, this process is often time-consuming and lacks the flexibility to dynamically adapt to varying content.
Recent works [17, 11, 31] leverage the powerful capabilities of pre-trained text-to-image diffusion models [40] to advance the design process. Despite significantly improving productivity and diversity, it remains challenging to flexibly portraying the visual characteristics (e.g., the transformation of texture and geometry) of input words without compromising readability. FontStudio [31] focuses solely on depicting the given word through style and texture while neglecting geometric deformation. In contrast, Word-as-image [17] and Khattat [16] target black-and-white designs by altering the shape of selected letters. WordArt Designer [11] attempts to integrate various transformations into semantic typography but may sacrifice readability when illustrating concepts that involve visible changes to the overall glyph geometry. For example, in Fig.5, the “tomato” resembles several tomatoes more than the character itself. MetaDesigner[12], as an integrated system employing multiple models, requires substantial manual effort and extensive engineering techniques to achieve promising artistic typography. Moreover, existing methods can express only a single interpretation of a given concept, further limiting flexibility and user control over the technique.
@@
We implement VitaGlyph through a three-phase approach: Knowledge Acquisition, Regional Decomposition, and Typography Stylization. In the first phase, we leverage large language models (e.g., ChatGPT [32]) to acquire prior knowledge of the input character, designing text prompts for the Subject and Surrounding. In the second phase, Grounding-DINO [23] is used to identify the part of the glyph that structurally resembles the subject concept, allowing the glyph image to be decomposed into subject and surrounding regions.
During the third phase, we render the subject and surrounding regions with their corresponding prompts through Controllable Compositional Generation, which we propose inventively. To better align with the intrinsic semantics of the subject concept, Semantic Typography is used to deform the geometry of the subject image before the rendering process. Notably, VitaGlyph enables users to incorporate multiple customized concepts into a single glyph, thereby enhancing user control and fostering greater creativity.
@@
To conclude, we summarize our main contributions as follows:
  • We introduce the training-free VitaGlyph, which enables flexible and imaginable artistic typography through controllable transformations of glyph geometry, while maintaining readability by dividing input characters into Subject and Surrounding.
  • VitaGlyph portrays the Subject as the essential concept of the input character and uses the surrounding to enrich the background without harming readability.
  • Comprehensive experiments on collected glyph images demonstrate that VitaGlyph surpasses competitors in both creativity and readability.
@@ 2.1 Artistic Text Generation
Artistic text generation [3] is divided into text stylization, which focuses on visual effects [44, 27], and Semantic Typography, which deforms text to match its meaning [42, 11, 16]. Models like eDiff-I[4], DeepFloyd [19], and TextDiffuser-2 [7] enhance generation using LLMs, while others [46, 26, 37] add motion.
However, these methods often capture only one layer of meaning. We propose a dual-branch model that combines stylization and typography for more complex text generation.
2.2 Large Language Models for Image Generation
Large language Models(LLMs) [1] have extend their impressive abilities from language tasks to vision synthesis. Most generation-related works [21, 47, 11] use LLMs for prompt understanding, layout planning and typography design. Due to these insights, we incorporate the LLM to design artistic characters.
@@ 2.3 Controllable and Compositional Generation
ControlNet-based methods [30, 49, 15, 24, 51] improve image synthesis control with conditions like segmentation or depth maps, and some studies [34, 48, 31] apply these to text generation. Leveraging the geometric nature of glyphs, we use ControlNet with glyph conditions as our foundation.
For multi-concept synthesis, works [6, 2, 29, 39] optimize noise maps for concept co-occurrence, while [18] uses Diffusion Models with LoRAs [14], and [25] integrates residual embeddings and layout priors.
Few methods [50] combine compositional generation with controllable conditions. Our work achieves pixel-level control in compositional generation using ControlNet and user-specific masks.
@@ 3 Method
Artistic typography generation task aims to visually convey the meaning of input characters in an appealing and comprehensible way while also ensuring legibility. To this end, we introduce a dual-branch method, namely VitaGlyph, which performs adaptive rendering on the Subject and Surrounding components of input glyph. The Subject undergoes structural modifications to flexibly express the intrinsic semantics of the input character, reflecting strong artistic essence. Meanwhile, the Surrounding enriches the details of the subject while maintaining overall readability by adhering to the character’s structure throughout the generation process.
As illustrated in Fig. 2, VitaGlyph consists of three phases: Knowledge Acquisition, Regional Decomposition, and Typography Stylization. In the Knowledge Acquisition phase, we ask the large language models (e.g., ChatGPT [32]) to acquire the prior knowledge of the Subject and Surrounding and design their text prompts. In Regional Decomposition phase, we parameterize the input character into glyph image, followed by employing Grounding-DINO (i.e., GDINO) to detect the subject part. Finally, during Typography Stylization phase, we firstly transform the subject’s structure and then render subject and surrounding images with their corresponding prompts through Controllable Compositional Generation (CCG).
@@ 3.1 Knowledge Acquisition
The goal of Knowledge Acquisition is to convert input characters into descriptive language prompts for subjects and surroundings that are interpretable by text-to-image diffusion models. Manually designing these prompts or relying solely on the input characters themselves (e.g., “rose”) can be challenging, as these methods are either time-consuming or produce descriptions that are difficult to interpret, particularly for abstract words like “win” or “lose” [11].
The large language models like ChatGPT [32] have demonstrated significant capabilities in understanding, imagining, and interpreting. Following prior work [11], we utilize ChatGPT [32] to generate concrete and detailed text descriptions for input characters. However, while [11] primarily focuses on depicting the basic concept of an input character, our approach aims to visualize it more completely and vividly as a compositioanl scene (i.e., subject and its surroundings). Specifically, given a query prompt $Q_{c}$ such as “Help me design an artistic character for …”, we employ ChatGPT to interpret the input character $c$ into a subject prompt $P_{\mathrm{sub}}$ and a surrounding prompt $P_{\mathrm{surr}}$ as follows:
@@
3.2 Regional Decomposition
Given subject description $P_{\mathrm{sub}}$, the Regional Decomposition phase aims to convert the input character $c$ into glyph image $I$ and then divide the image $I$ into subject and surrounding regions at the image level. The subject region corresponds to the part of the image that is most structurally similar to the essential concept described in $P_{\mathrm{sub}}$.
3.2.1 Parameterization
Given an input character $c$, we use the FreeType font library [8] to extract the outline of the character and convert it into cubic Bézier curves. After that, the Differentiable Vector Graphics(DiffVG) [20] is utilized to raster obtained Bézier curves into get a final glyph image $I$.
3.2.2 Region Interpretation
@@
To preserve the shape of input glyph (i.e., readability) as much as possible, we choose the region most similar to the structure of subject concept as subject region (e.g., “snow” in Fig. 3). Compared to previous works that use random region selection [11, 12] or manual region selection [17], we leverage Grounding-DINO [23] to detect the part of input image $I$ that matches the subject description $P_{\mathrm{sub}}$, thereby efficiently bridging the semantic gap between the selected regions and the subjects, while significantly reducing time. As the detected results contain multiple bounding boxes of varying sizes, we filter out those with unreasonable sizes and any that fall below the confidence threshold [^1] We set the reasonable proportion of the bounding box area relative to the image area between 0.4 and 0.6, and the confidence threshold to 0.5 Then we select the bounding box with the highest confidence score as the subject region:
$I_{\mathrm{sub}},M=\texttt{GDINO}_{\mathrm{rank}}(I,P_{\mathrm{sub}}),$ (2)
where $I_{\mathrm{sub}}$, $\texttt{GDINO}_{\mathrm{rank}}$ and $M$ denote the subject image, Grounding-DINO with filtering and ranking strategies and the mask area of the Subject, respectively. As shown in Fig. 3, due to Grounding-DINO’s open-domain detection capabilities, the $I_{\mathrm{sub}}$(red bounding box) of characters that best resemble the $P_{\mathrm{sub}}$ can be easily detected. There are no specific structural requirements for the surrounding image $I_{\mathrm{surr}}$, which is expected to maintain legibility. Thus, we consider the remaining part of the glyph image as the surrounding region:
$I_{\mathrm{surr}}=I-I_{\mathrm{sub}}.$ (3)
3.3 Typography Stylization
As mentioned above, the subject and surrounding components play different roles in depicting the visual semantics of input characters, so their structures and textures are rendered in different ways. Firstly, the structure of subject image is flexibly altered and enriched by Semantic Typography, while the surrounding remains unchanged. Secondly, we render their texture style through two independent ControlNets (i.e., as $\mathrm{ControlNet}_{\mathrm{sub}}$ and $\mathrm{ControlNet}_{\mathrm{surr}}$) in a compositional manner, which we term as Controllable Compositional Generation (CCG).
3.3.1 Semantic Typography
Previous works [17, 11] rely on time-consuming score distillation sampling loss [36] to transform the structure of input glyphs. To improve efficiency, we directly utilize the depth-to-image model in LDM [40] to adjust the geometry of $I_{\mathrm{sub}}$. Specifically, we adopt the SDEdit algorithm [28] to perform geometry transformation, conditioned on the depth map $I_{\mathrm{sub}}$ and subject prompt $P_{\mathrm{sub}}$. It can be described as follows:
@@ 3.3.1 Semantic Typography
3.3.2 Controllable Compositional Generation
Figure 4: Overview of Controllable Compositional Generation. The Controllable Compositional Generation utilizes Stable Diffusion and two independent ControlNets to synthesize and fuse the subject and surrounding regions with the given mask $M$.
Figure 5: Visualization of artistic typography results compared with existing methods. WordArt Designer and MetaDesigner apply transformations to entire characters, often sacrificing readability, as seen in “mushroom”, where the bottom becomes unrecognizable. Additionally, some characters blend into the background without independent diffusion modeling, such as “cabinet” in WordArt and “tomato” in MetaDesigner. In contrast, our method preserves better readability with controllable transformations, separates characters from the background and maintains a visually engaging design. We also provide a comparison with Word-as-image in the English version.
To render subject and surrounding images with their corresponding prompts independently, we apply two ControlNets to separately transfer texture style in a compositional manner. For the subject image $I_{\mathrm{sub}}$, after transforming its structure in SemTypo into $I^{*}_{\mathrm{sub}}$, we use the segmentation version of ControlNet, referred to as $\mathrm{ControlNet}_{\mathrm{sub}}$. For the surrounding image $I_{\mathrm{surr}}$, we select the scribble version of ControlNet, denoted as $\mathrm{ControlNet}_{\mathrm{surr}}$ to maintain the character layout while adding creativity and artistic flair. This dual-branch diffusion process decouples the rendering of subject and surrounding, allowing the generated artistic typography to incorporate different concepts.
Specifically, we first encode $P_{\mathrm{sub}}$ and $P_{\mathrm{surr}}$ into text embeddings $\boldsymbol{p}_{\mathrm{sub}}$ and $\boldsymbol{p}_{\mathrm{surr}}$, respectively. Both $\mathrm{ControlNet}_{\mathrm{sub}}$ and $\mathrm{ControlNet}_{\mathrm{surr}}$ share the same random noise $\epsilon_{T}\in\mathcal{N}(0,I)$ as the initial latent, conditioned on their corresponding structures and text embeddings. We use the DDIM algorithm [41] for sampling. To fuse subject and surrounding textures at timestep $t$, given subject mask $M$, we combine the predicted noises $\epsilon^{\mathrm{sub}}_{t}$ and $\epsilon^{\mathrm{surr}}_{t}$ to compute the overall noise $\epsilon^{\mathrm{overall}}_{t}$ as:

### v3

@@ Abstract
Artistic typography is a technique to visualize the meaning of input character in an imaginable and readable manner. With powerful text-to-image diffusion models, existing methods directly design the overall geometry and texture of input character, making it challenging to ensure both creativity and legibility. In this paper, we introduce a dual-branch, training-free method called VitaGlyph, enabling flexible artistic typography with controllable geometry changes while maintaining the readability. The key insight of VitaGlyph is to treat input character as a scene composed of a Subject and its Surrounding, which are rendered with varying degrees of geometric transformation. To enhance the visual appeal and creativity of the generated artistic typography, the subject flexibly expresses the essential concept of the input character, while the surrounding enriches relevant background without altering the shape, thus maintaining overall readability. Specifically, we implement VitaGlyph through a three-phase framework: (i) Knowledge Acquisition leverages large language models to design text descriptions for the subject and surrounding. (ii) Regional Interpretation detects the part that most closely matches the subject description and refines the structure via Semantic Typography. (iii) Attentional Compositional Generation separately renders the textures of the Subject and Surrounding regions and blends them in an attention-based manner. Experimental results demonstrate that VitaGlyph not only achieves better artistry and readability but also manages to depict multiple customized concepts, facilitating more creative and pleasing artistic typography generation. Our code will be made publicly available.
Figure 1: Examples of our VitaGlyph. The rendered images not only convey the intrinsic semantics of input characters (first row) but also visualize the user-specific concepts (second row). Sub. denotes the deformable glyph region that conveys the semantics; Surr. denotes the texture in remaining area.
@@ 1 Introduction
Figure 2: Overview of VitaGlyph. VitaGlyph divides the input glyph image into Subject and Surrounding and follows a three-phase pipeline: (i) Given an input character, Knowledge Acquisition leverages LLMs to design prompts for both subject and surrounding. (ii) With the subject prompt and glyph image, Regional Interpretation detects the part that best matches the subject description and transform the subject’s structure via Semantic Typography. (iii) Attentional Compositional Generation separately renders textures for both regions through attention-based mechanism.
Artistic typography focuses on emphasizing the visual meaning of words while maintaining their legibility. This technique has widespread applications across various fields such as commercial advertising (Gao et al. 2023; Lin et al. 2023), social education (Vungthong, Djonov, and Torr 2017; Duda et al. 2023), and design promotion (Zhao and Lian 2023; Xiao et al. 2024). Traditional artistic typography involves manually applying design principles such as color, spacing, and font choice to convey meaning (He et al. 2023; Iluz et al. 2023). However, this process is often time-consuming and lacks the flexibility to dynamically adapt to varying contents and concepts.
Recent works (Iluz et al. 2023; He et al. 2023; Mu et al. 2024) leverage the powerful capabilities of pre-trained text-to-image diffusion models (Rombach et al. 2022) to advance the design process. Despite significantly improving productivity and diversity, it remains challenging to flexibly portray the visual characteristics (e.g., the transformation of texture and geometry) of input words while maintaining good readability. FontStudio (Mu et al. 2024) focuses solely on depicting the given word through style and texture while neglecting geometric deformation. In contrast, Word-as-image (Iluz et al. 2023) and Khattat (Hussein et al. 2024) target black-and-white designs by altering the shape of selected letters. WordArt Designer (He et al. 2023) applies distorted transformations to the entire word when attempting to convey multiple concepts, which may sacrifice readability due to significant changes in the overall glyph structure. For example, in Fig. 5, the “tomato” resembles several tomatoes more than the character itself. MetaDesigner(He et al. 2024), as an integrated system that employs multiple models, requires substantial manual effort and extensive engineering techniques to achieve promising artistic typography. Moreover, current methods are constrained by their ability to convey only a singular interpretation of a given concept, thereby further restricting flexibility and limiting users’ control in the creation of artistic typography.
@@
We implement VitaGlyph through a three-phase approach: Knowledge Acquisition, Regional Interpretation, and Typography Stylization. In the first phase, we leverage large language models (e.g., ChatGPT (OpenAI 2024a)) to acquire prior knowledge of the input character, designing text prompts for the Subject and Surrounding. In the second phase, different from previous works, we utilize Grounding-DINO (Liu et al. 2023a) to identify the part of the input glyph that most resembles the subject concept. It allows the decomposed subject inherently contain semantic priors, with which Semantic Typography is used to deform the geometry of the subject image naturally. During the third phase, we render the subject/surrounding regions with their corresponding prompts through Attentional Compositional Generation(ACG). To better enhance the aesthetic qualities, we inventively propose Cross-branch Attention and Attention-driven Control Fusion.
@@
We summarize our main contributions as follows:
  • We introduce VitaGlyph, a region‐aware framework for flexible, expressive artistic typography via independent stylization without sacrificing legibility.
  • We formalize the dual concepts of Subject and surrounding. Subject transforms a glyph’s core structure for artistic expression and surrounding only enriches decorative details for preserving readability.
  • VitaGlyph integrates Attention‐Guided Compositional Generation, enhancing aesthetic quality and producing more natural blending.
@@ 2.1 Artistic Text Generation
Artistic text generation (Bai et al. 2024) is divided into text stylization, which focuses on visual effects (Wang et al. 2023; Mao et al. 2022), and Semantic Typography, which deforms text to match its meaning (Tanveer et al. 2023; He et al. 2023; Hussein et al. 2024). Models like eDiff-I(Balaji et al. 2022), DeepFloyd (Lab 2023), and TextDiffuser-2 (Chen et al. 2023) enhance generation using LLMs, while others (Xie et al. 2023; Liu et al. 2024b; Pu et al. 2024) add motion. However, these methods often capture only one layer of meaning. We propose a dual-branch model that combines stylization and typography for more complex text generation.
2.2 LLMs for Image Generation
Large language Models(LLMs) (Achiam et al. 2023) have extend their impressive abilities from language tasks to vision synthesis. Most generation-related works (Lian et al. 2023; Yang et al. 2024a; He et al. 2023) use LLMs for prompt understanding, layout planning and typography design. Due to these insights, we incorporate the LLM to design artistic characters.
@@ 2.3 Controllable and Compositional Generation
ControlNet-based methods (Mou et al. 2024; Zhang, Rao, and Agrawala 2023; Huang et al. 2023; Liu et al. 2024a; Zhang et al. 2023) improve image synthesis control with conditions like segmentation or depth maps, and some studies (Peong, Uchida, and Haraguchi 2024; Yang et al. 2024b; Mu et al. 2024) apply these to text generation. Leveraging the geometric nature of glyphs, we use ControlNet with glyph conditions as our foundation.
For multi-concept synthesis, works (Chefer et al. 2023; Agarwal et al. 2023; Meral et al. 2024; Rassin et al. 2024) optimize noise maps for concept co-occurrence, while (Jiang et al. 2024) uses Diffusion Models with LoRAs (Hu et al. 2021), and (Liu et al. 2023b) integrates residual embeddings and layout priors.
Few methods (Zhang et al. 2024) combine compositional generation with controllable conditions. Our work achieves pixel-level control in compositional generation using ControlNet and user-specific masks.
@@ 3 Method
Artistic typography generation task aims to visually convey the meaning of input characters in an appealing and comprehensible way while also ensuring legibility. To this end, we introduce a dual-branch method, namely VitaGlyph, which performs adaptive rendering on the Subject and Surrounding components of input glyph. The Subject undergoes structural modifications to flexibly express the intrinsic semantics of the input character, reflecting a strong artistic essence. Meanwhile, the Surrounding enriches the details of the subject while maintaining overall readability by adhering to the character’s structure throughout the generation process.
As illustrated in Fig. 2, VitaGlyph consists of three phases: Knowledge Acquisition, Regional Interpretation, and Attentional Compositional Generation. In the Knowledge Acquisition phase, we ask large language models (e.g., ChatGPT) to acquire the prior knowledge of the Subject and Surrounding and design their text prompts. In Regional Interpretation phase, we parameterize the input character into a glyph image, followed by employing Grounding-DINO (i.e., GDINO) to detect the subject part and Semantic Typography (i.e., SemTypo) to transform the structure. Finally, during Attentional Compositional Generation (ACG) phase, we render subject and surrounding images with their corresponding prompts through Cross-branch Attention blending and Attention-driven Control Fusion.
@@ 3.1 Knowledge Acquisition
The goal of Knowledge Acquisition is to convert input characters into descriptive language prompts for subjects and surroundings that are interpretable by text-to-image diffusion models. Manually designing these prompts or relying solely on the input characters themselves (e.g., “rose”) can be challenging, as these methods are either time-consuming or produce descriptions that are difficult to interpret, particularly for abstract words like “win” or “lose” (He et al. 2023).
large language models such as ChatGPT (OpenAI 2024a) have demonstrated significant capabilities in understanding, imagining, and interpreting. We use ChatGPT (OpenAI 2024a) to generate detailed and concrete text descriptions for input characters. However, while related work using LLM focuses primarily on depicting the basic concept of an input character, our approach aims to visualize it more completely and vividly as a compositional scene (i.e., subject and its surroundings). Specifically, given a query prompt $Q_{c}$ such as “Help me design an artistic character for…”, we employ ChatGPT to interpret the input character $c$ into a subject prompt $P_{\mathrm{sub}}$ and a surrounding prompt $P_{\mathrm{surr}}$ as follows:
@@
3.2 Regional Interpretation
@@
The Regional Interpretation phase aims to convert the input character $c$ into glyph image $I$ and then divide the image $I$ into subject and surrounding regions at the image level firstly. The subject region corresponds to the part of the image that is most structurally similar to the essential concept described in $P_{\mathrm{sub}}$. Then it will undergo structural modifications to flexibly express the intrinsic semantics through Semantic Typography (SemTypo).
Region Selection
Given an input character $c$, we use ImageFont and ImageDraw of the PIL library to rasterize it into a glyph image $I$ based on a predefined TrueType font.
To preserve the shape of input glyph (i.e., readability) as much as possible, we choose the region most similar to the structure of subject concept as subject region (e.g., “snow” in Fig. 3). Compared to previous works that use random region selection (He et al. 2023, 2024) or manual region selection (Iluz et al. 2023), we leverage Grounding-DINO (Liu et al. 2023a) to detect the part of input image $I$ that matches the subject description $P_{\mathrm{sub}}$. Since Grounding-DINO encapsulates rich visual-semantic knowledge, the detected regions inherently contain semantic priors in terms of font structure, thereby efficiently bridging the semantic gap between the selected regions and the subjects, while significantly reducing time.
As the detected results contain multiple bounding boxes of varying sizes, we filter out those with unreasonable sizes and any that fall below the confidence threshold [^1] We set the reasonable proportion of the bounding box area relative to the image area between 0.4 and 0.6, and the confidence threshold to 0.5 Then we select the bounding box with the highest confidence score as the subject region and the remaining of the glyph image as surrounding image:
$I_{\mathrm{sub}},M=\texttt{GDINO}_{\mathrm{rank}}(I,P_{\mathrm{sub}}),I_{\mathrm{surr}}=I-I_{\mathrm{sub}},$ (2)
where $I_{\mathrm{sub}}$, $\texttt{GDINO}_{\mathrm{rank}}$, $M$ and $I_{\mathrm{surr}}$ denote the subject image, Grounding-DINO with the filtering and ranking strategy, the mask area and the surrounding image, respectively. As shown in Fig. 3 and Fig. 7, due to Grounding-DINO’s open-domain detection capabilities, $I_{\mathrm{sub}}$ (red bounding box) of characters that best resemble $P_{\mathrm{sub}}$ can be easily detected.
Semantic Typography
Previous works (Iluz et al. 2023; He et al. 2023) rely on the time-consuming Score Distillation Sampling loss (Poole et al. 2022) to transform the structure of input glyphs. Based on the detected $I_{\mathrm{sub}}$, we directly use the Depth-to-image model in LDM (Rombach et al. 2022) to adjust the geometry of $I_{\mathrm{sub}}$, improving efficiency simultaneously. Specifically, we adopt the SDEdit algorithm (Meng et al. 2021) to perform geometry transformation, conditioned on the depth map $I_{\mathrm{sub}}$ and the subject prompt $P_{\mathrm{sub}}$. It can be described as follows:
@@
3.3 Attentional Compositional Generation
Figure 4: Overview of Attentional Compositional Generation. The ACG module utilizes Stable Diffusion and two independent ControlNets to synthesize and fuse the subject and surrounding regions with the given mask $M$.
As mentioned above, the Subject and surrounding play different roles in depicting the visual semantics of input characters, so their structures and textures are rendered in different ways through two independent ControlNets. Rather than simple noise‐based blending, we compose them in an attention-based compositional manner termed as Attentional Compositional Generation (ACG). It involves Cross-branch Attention mechanism and Attention-driven Control Fusion module.
Figure 5: Visualization of artistic typography results compared with existing methods. WordArt and MetaDesigner apply transformations to entire characters, often sacrificing readability, as seen in “mushroom”, where the bottom becomes unrecognizable. Besides, some characters blend into the background without independent diffusion modeling, such as “cabinet” in WordArt and “tomato” in MetaDesigner. Additionally, former works lack some key semantics (i.e., ”Earth” in WordArt and “robot” in Metadesigner in “Hello World”). In contrast, our method preserves better readability with controllable transformations, separates characters from the background and maintains a visually engaging design.
To render subject and surrounding images with their corresponding prompts independently, we apply two ControlNets to separately transfer texture style in a compositional manner. For the subject image $I_{\mathrm{sub}}$, after transforming its structure in SemTypo into $I^{*}_{\mathrm{sub}}$, we use the segmentation version of ControlNet, termed as $\mathrm{ControlNet}_{\mathrm{sub}}$. Due to the fact that the subject region has semantic priors structurally, the subsequent deformation of this region is more naturally and less likely to suffer significant distortion than that of previous works. For the surrounding image $I_{\mathrm{surr}}$, we select the scribble version of ControlNet, denoted as $\mathrm{ControlNet}_{\mathrm{surr}}$ to maintain the character layout while adding creativity and artistic flair. This dual-branch diffusion process decouples the rendering of subject and surrounding, allowing the generated artistic typography to incorporate different concepts.
Specifically, we first encode $P_{\mathrm{sub}}$ and $P_{\mathrm{surr}}$ into text embeddings $\boldsymbol{p}_{\mathrm{sub}}$ and $\boldsymbol{p}_{\mathrm{surr}}$, respectively. Both $\mathrm{ControlNet}_{\mathrm{sub}}$ and $\mathrm{ControlNet}_{\mathrm{surr}}$ share the same random noise $\epsilon_{T}\in\mathcal{N}(0,I)$ as the initial latent, conditioned on their corresponding structures and text embeddings. We use the DDIM algorithm (Song, Meng, and Ermon 2020) for sampling. To fuse subject and surrounding textures at timestep $t$, given subject mask $M$, we combine the predicted noises $\epsilon^{\mathrm{sub}}_{t}$ and $\epsilon^{\mathrm{surr}}_{t}$ to compute the overall noise $\epsilon^{\mathrm{overall}}_{t}$ as: